# Convergence of a Finite Volume Scheme for Compactly Heterogeneous Scalar Conservation Laws

### v2

@@ 3.2 Steady states and stability
Assume that $H$ satisfies (C3)–(CVX). Then, there exists a unique function $\alpha\in\mathbf{C}^{2}(\mathbb{R},\mathbb{R})$ such that for all $x\in\mathbb{R}$, $\partial_{u}H(x,\alpha(x))=0$. Moreover,
$\forall x\in\mathbb{R},\quad|x|\geq X\implies\alpha^{\prime}(x)=0.$
@@
Introduce $J\in\mathbb{N}^{*}$ such that $X\in I_{J}$. Recall that the space dependency of $H$ is localized in $[-X,X]$.
@@
$j\geq J+1\implies\alpha_{j}=\alpha(X)\quad\text{and}\quad j\leq-J-1\implies\alpha_{j}=\alpha(-X).$
@@
It is worth mentioning that for all $j\in\mathbb{Z}$, $F_{\mathrm{int}}^{j+1/2}$ is a monotone, locally Lipschitz, numerical flux, see [2, Section 4]. Also, let us precise that in general, $F_{\mathrm{int}}^{j+1/2}$ is not consistent, meaning that for all $k\in\mathbb{R}$, $F_{\mathrm{int}}^{j+1/2}(k,k)$ is not equal to $h_{j}(k)$ or to $h_{j+1}(k)$. However, using the regularity of $H$, we can prove that a consistency holds, at the limit.
@@
Assume that $H$ satisfies (C3)–(CVX). Then for any $k\in\mathbb{R}$, there exists a positive constant $\mu_{k}$ such that for all $j\in\mathbb{Z}$,
@@
$F_{\mathrm{int}}^{j+1/2}(k,k)=\left\{\begin{array}[]{cclc}\max\{h_{j}(\alpha_{j}),h_{j+1}(k)\}&\text{if}&k\leq\alpha_{j},\alpha_{j+1}&\text{(I)}\\ \max\{h_{j}(\alpha_{j}),h_{j+1}(\alpha_{j+1})\}&\text{if}&\alpha_{j+1}\leq k\leq\alpha_{j}&\text{(II)}\\ \max\{h_{j}(k),h_{j+1}(k)\}&\text{if}&\alpha_{j}\leq k\leq\alpha_{j+1}&\text{(III)}\\ \max\{h_{j}(k),h_{j+1}(\alpha_{j+1})\}&\text{if}&k>\alpha_{j},\alpha_{j+1}&\text{(IV)}\end{array}\right.$
The proof reduces to a case by case study. Set $\delta\coloneq\left|F_{\mathrm{int}}^{j+1/2}(k,k)-h_{j+1}(k)\right|$
@@
$\delta=h_{j}(\alpha_{j})-h_{j+1}(k)\leq h_{j}(k)-h_{j+1}(k)\leq\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x.$
@@
$|h_{j}(\alpha_{j})-h_{j+1}(k)|\leq\left(\|\alpha^{\prime}\|_{\mathbf{L}^{\infty}}\sup_{\underset{|p|\leq\|\alpha\|_{\mathbf{L}^{\infty}}}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|+\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\right)\Delta x.$
Case (III) In that case, we can bound $\delta$ by $\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x$.
Case (IV) If $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j}(k)$, then we bound $\delta$ by $\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x$. Otherwise, $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j+1}(\alpha_{j+1})$, meaning that $h_{j+1}(\alpha_{j+1})\geq h_{j}(k)$. Therefore,
$\delta$ $\leq|h_{j+1}(\alpha_{j+1})-h_{j}(k)|+|h_{j}(k)-h_{j+1}(k)|$
$=(h_{j+1}(\alpha_{j+1})-h_{j}(k))+|h_{j}(k)-h_{j+1}(k)|$
$\leq 2|h_{j}(k)-h_{j+1}(k)|$
$\leq 2\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x.$
@@
To prove (M1), we explicitly construct steady states of the scheme (3.12).
@@
Following Section 3.1, it will be useful to introduce for all $j\in\mathbb{Z}$ and $y\in\mathopen{]}\min h_{j},+\infty\mathclose{[}$, the two solutions $S_{j}^{-}(y)<\alpha_{j}<S_{j}^{+}(y)$ of the equation $h_{j}(x)=y$, $x\in\mathbb{R}$.
@@
Assume that $H$ satisfies (C3)–(CVX). Let $c\in\mathbb{R}$ such that $c\geq\sup_{\mathbb{R}}\alpha$. Define for all $j\in\mathbb{Z}$,
@@
Proof.  The key point is that for all $j\in\mathbb{Z}$, $(v_{j},v_{j+1})$ and $(w_{j},w_{j+1})$ are chosen in the germ $\mathcal{G}_{j+1/2}$, see Definition 3.1, which ensures by Proposition 3.3 that
@@
$v_{j+1}=\left\{\begin{array}[]{ccl}c&\text{if}&j\leq-J-2\\ S_{j+1}^{-}(h_{j}(v_{j}))&\text{if}&j\geq-J-1,\end{array}\right.\quad\quad w_{j}=\left\{\begin{array}[]{ccl}S_{j}^{-}(h_{j+1}(w_{j}))&\text{if}&j\leq J\\ c&\text{if}&j\geq J+1.\end{array}\right.$
@@
Assume that $H$ satisfies (C3)–(CVX). Let $(m,M)\in\mathbb{R}^{2}$ such that $m\leq\inf_{\mathbb{R}}\alpha$ and $M\geq\sup_{\mathbb{R}}\alpha$. Then there exist two steady states of (3.12), $(m_{j})_{j}$ and $(M_{j})_{j}$, such that:
(i) $(m_{j})_{j}$ is bounded by below ;
(ii) $(M_{j})_{j}$ is bounded by above ;
(iii) for all $j\in\mathbb{Z}$, $m_{j}\leq m$ and $M_{j}\geq M$.
Proof.  We only give the details for the construction of $(M_{j})_{j}$.
Set $\mathcal{U}_{h}\coloneq\max_{x\in\mathbb{R}}H(x,M)$. By monotonicity, for all $j\in\mathbb{Z}$, $S_{j}^{+}(\mathcal{U}_{h})\geq M$. Therefore, if $\overline{M}\coloneq\max_{j\in\mathbb{Z}}S_{j}^{+}(\mathcal{U}_{h})$, then $\overline{M}\geq M\geq\sup_{x\in\mathbb{R}}\alpha(x)$. Define
$\forall j\in\mathbb{Z},\quad M_{j+1}\coloneq\left\{\begin{array}[]{ccl}\overline{M}&\text{if}&j\leq-J-2\\ S_{j+1}^{+}(h_{j}(M_{j}))&\text{if}&j\geq-J-1.\end{array}\right.$
Clearly, $(M_{j})_{j}$ is a steady state of the scheme.
@@
for all $j\in\mathbb{Z}$, $M_{j}\geq M$. Indeed, for all $j\in\mathbb{Z}$,
$h_{j}(M_{j})=h_{j+1}(M_{j+1})=h_{-J-1}(\max_{i\in\mathbb{Z}}S_{i}^{+}(\mathcal{U}_{h}))\geq h_{-J-1}(S_{-J-1}^{+}(\mathcal{U}_{h}))=\mathcal{U}_{h}\geq h_{j}(M),$
from which we deduce, by monotonicity, that $M_{j}\geq M$. The claim is proved.
@@
$(M_{j})_{j}$ is bounded by above. By definition, for all $j\in\mathbb{Z}$,
$M_{j}=S_{j}^{+}\left(H(-X,\overline{M})\right)\leq\max_{x\in\mathbb{R}}h_{+}^{-1}\left(x,H(-X,\overline{M})\right)\coloneq\overline{u},$
where for all $x\in\mathbb{R}$, we denote by $u\mapsto h_{+}^{-1}(x,u)$ the reciprocal of $u\mapsto h(x,u)$ on $[\alpha(x),+\infty\mathclose{[}$. We see that $\overline{u}$ only depends on $H$ and $M$.   $\square$
Proof of (M1).  With reference to Lemma 3.6, fix $(m,M)\in\mathbb{R}^{2}$ satisfying $m\leq\inf_{\mathbb{R}}\alpha$ and $M\geq\sup_{\mathbb{R}}\alpha$, such that for a.e. $x\in\mathbb{R}$, $m\leq u_{o}(x)\leq M$. Consider $(m_{j})_{j}$ and $(M_{j})_{j}$, steady states given by Lemma 3.10.
@@
$(m_{\Delta},M_{\Delta})\coloneq\sum_{j\in\mathbb{Z}}(m_{j},M_{j})\mathbf{1}_{I_{j}},$
and $\underline{u}$, $\overline{u}$ as they were in the proof of Lemma 3.10.
The key point is that $m_{\Delta}$ and $M_{\Delta}$ are both stationary solutions to
$\partial_{t}\mathcal{U}(t,x)+\partial_{x}\bigl{(}H(w_{\Delta}(x),\mathcal{U}(t,x))\bigr{)}=0.$
Therefore, for all $n\in\mathbb{N}$, (D2) of Theorem 3.5 ensures that
$m_{\Delta}\leq\rho^{n}\leq M_{\Delta}\implies\forall t\in[t^{n},t^{n+1}],\quad m_{\Delta}\leq\mathcal{U}^{n}(t)\leq M_{\Delta}.$
Since by construction of the sequences $(m_{j})_{j\in\mathbb{Z}}$ and $(M_{j})_{j\in\mathbb{Z}}$, we have $m_{\Delta}\leq\rho^{0}\leq M_{\Delta}$, a direct induction provides:
$\forall t\in\mathbb{R}^{+},\quad m_{\Delta}\leq u_{\Delta}(t)\leq M_{\Delta},$
immediately leading to (2.4) by definition of $\underline{u}$ and $\overline{u}$.
The proof is concluded.   $\square$
@@ 3.3 Convergence of the scheme
We first establish the strong compactness of $(u_{\Delta})_{\Delta}$ with a compensated compactness argument.

### v3

@@ 3.2 Steady states and stability
Assume that $H$ satisfies (C3)–(CH)–(CVX). Then, there exists a unique function $\alpha\in\mathbf{C}^{2}(\mathbb{R},\mathbb{R})$ such that for all $x\in\mathbb{R}$, $\partial_{u}H(x,\alpha(x))=0$. Moreover,
$\forall x\in\mathbb{R},\quad{\left|x\right|}\geq X\implies\alpha^{\prime}(x)=0.$
@@
Introduce $J\in\mathbb{N}^{*}$ such that $X\in I_{J}$. Recall that the space dependency of $H$ is localized in [X,X][-X,X].
@@
$(j\geq J+1\implies\alpha_{j}=\alpha(X))\quad\text{and}\quad(j\leq-J-1\implies\alpha_{j}=\alpha(-X)).$
@@
It is worth mentioning that for all $j\in\mathbb{Z}$, $F_{\mathrm{int}}^{j+1/2}$ is a monotone, locally Lipschitz, numerical flux, see [AJG2004, Section 4]. Also, remark that in general, $F_{\mathrm{int}}^{j+1/2}$ is not consistent, meaning that for all $k\in\mathbb{R}$, $F_{\mathrm{int}}^{j+1/2}(k,k)$ is not equal to $h_{j}(k)$ or to $h_{j+1}(k)$. However, using the regularity of $H$, we can prove that a consistency holds, at the limit.
@@
Assume that $H$ satisfies (C3)–(CH)–(CVX). Then for any $k\in\mathbb{R}$, there exists a positive constant $\mu_{k}$ such that for all $j\in\mathbb{Z}$,
@@
$F_{\mathrm{int}}^{j+1/2}(k,k)=\left\{\begin{array}[]{cclc}\max\{h_{j}(\alpha_{j}),h_{j+1}(k)\}&\text{if}&k\leq\alpha_{j},\alpha_{j+1}\quad&\text{(I)}\\ \max\{h_{j}(\alpha_{j}),h_{j+1}(\alpha_{j+1})\}&\text{if}&\alpha_{j+1}\leq k\leq\alpha_{j}\quad&\text{(II)}\\ \max\{h_{j}(k),h_{j+1}(k)\}&\text{if}&\alpha_{j}\leq k\leq\alpha_{j+1}\quad&\text{(III)}\\ \max\{h_{j}(k),h_{j+1}(\alpha_{j+1})\}&\text{if}&k>\alpha_{j},\alpha_{j+1}\quad&\text{(IV)}\end{array}\right$.
The proof reduces to a case by case study. Set $\delta\coloneq{\left|F_{\mathrm{int}}^{j+1/2}(k,k)-h_{j+1}(k)\right|}$.
@@
$\delta=h_{j}(\alpha_{j})-h_{j+1}(k)\leq h_{j}(k)-h_{j+1}(k)\leq\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\;\Delta x.$
@@
${\left|h_{j}(\alpha_{j})-h_{j+1}(k)\right|}\leq\bigl({\left\|\alpha^{\prime}\right\|}_{\mathbf{L}^{\infty}(\mathbb{R})}\sup_{\underset{{\left|p\right|}\leq{\left\|\alpha\right\|}_{\mathbf{L}^{\infty}(\mathbb{R})}}{x\in\mathbb{R}}}{\left|\partial_{u}H(x,p)\right|}+\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\bigr)\Delta x.$
Case (III) In that case, we can bound $\delta$ by $\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\;\Delta x$.
Case (IV) If $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j}(k)$, then we bound $\delta$ by $\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\;\Delta x$. Otherwise, $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j+1}(\alpha_{j+1})$, meaning that $h_{j+1}(\alpha_{j+1})\geq h_{j}(k)$. Therefore,
$\delta$ $\leq{\left|h_{j+1}(\alpha_{j+1})-h_{j}(k)\right|}+{\left|h_{j}(k)-h_{j+1}(k)\right|}$
$=(h_{j+1}(\alpha_{j+1})-h_{j}(k))+{\left|h_{j}(k)-h_{j+1}(k)\right|}$
$\leq 2{\left|h_{j}(k)-h_{j+1}(k)\right|}$
$\leq 2\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\;\Delta x.$
@@
To prove the $\mathbf{L}^{\infty}$ bounds (M1), we explicitly construct steady states of the scheme (3.12).
@@
Following Section 3.1, it will be useful to introduce for all $j\in\mathbb{Z}$ and $y\in\mathopen{]}\min h_{j},+\infty\mathclose{[}$, the two solutions $S_{j}^{-}(y)<\alpha_{j}<S_{j}^{+}(y)$ of the equation $h_{j}(x)=y$, $x\in\mathbb{R}$. For all $j\in\mathbb{Z}$, the definition of the germ $\mathcal{G}_{j+1/2}$ follows from Definition 3.1.
@@
Assume (C3)–(CH)–(CVX) hold. Let $c\in\mathbb{R}$ such that $c\geq\sup_{\mathbb{R}}\alpha$. Define for all $j\in\mathbb{Z}$,
@@
Proof.  Consider $(v_{j})_{j}$. Since $c\geq\sup_{\mathbb{R}}\alpha$, for all $j\in\mathbb{Z}$, $(v_{j},v_{j+1})\in\mathcal{G}_{j+1/2}$, see Definition 3.1. Indeed, we have $v_{j}\geq\alpha_{j}$, $v_{j+1}\geq\alpha_{j+1}$ and $h_{j}(v_{j})=h_{j+1}(v_{j+1})$. Consequently, Proposition 3.3 ensures that
@@
$v_{j+1}=\left\{\begin{array}[]{ccl}c&\text{if}&j\leq-J-2\\ S_{j+1}^{-}(h_{j}(v_{j}))&\text{if}&j\geq-J-1,\end{array}\right.\quad\quad w_{j}=\left\{\begin{array}[]{ccl}S_{j}^{-}(h_{j+1}(w_{j+1}))&\text{if}&j\leq J\\ c&\text{if}&j\geq J+1.\end{array}\right$.
For the construction of the steady states, we will use geometric properties of both $H$ and its Legendre transform with respect to the second variable (see [CannarsaSinestrariSBook, Appendix 4.2]). We call it $L$ from now on.
@@
Assume that $H$ satisfies (C3)–(CH)–(CVX). Then the following points hold.
(i) $L\in\mathbf{C}^{2}(\mathbb{R}^{2},\mathbb{R})$ and satisfies (CH)–(CVX).
(ii) We have
$\forall\lambda>0,\;\forall x\in\mathbb{R},\;\forall p\in\mathbb{R},\quad\lambda{\left|p\right|}-H(x,p)\leq\sup_{\underset{{\left|v\right|}\leq\lambda}{y\in\mathbb{R}}}L(y,v).$ (3.13)
Proof.  Recall that $L(x,v)\coloneq\sup_{p\in\mathbb{R}}(pv-H(x,p))$, see [CannarsaSinestrariSBook, Appendix 4.2]. The growth of $H$ ensures that $L$ is well-defined on $\mathbb{R}^{2}$.
(i) The regularity of $L$ follows from an application of the Implicit Function Theorem and the fact that $L$ satisfies (CH) is immediate from the definition. Finally, for all $x\in\mathbb{R}$, $v\mapsto\partial_{v}L(x,v)$ is the reciprocal of $u\mapsto\partial_{u}H(x,u)$, therefore, (CVX) holds as well.
(ii) We first prove that (3.13) holds swapping the roles of $H$ and $L$.
For all $\lambda>0$, $x\in\mathbb{R}$ and $v\in\mathbb{R}\setminus\{0\}$, specializing with $p=\frac{\lambda v}{{\left|v\right|}}$ in the definition of $L$ yields:
$L(x,v)\geq\lambda{\left|v\right|}-H(x,\frac{\lambda v}{{\left|v\right|}})\geq\lambda{\left|v\right|}-\sup_{\underset{{\left|p\right|}\leq\lambda}{y\in\mathbb{R}}}H(y,p).$
Inequality (3.13) follows from the same computations, using the fact that the Legendre transform is an involution on convex functions.   $\square$
We are now in position to construct steady states of the scheme.
Lemma 3.11.
Assume that $H$ satisfies (C3)–(CH)–(CVX). Let $(m,M)\in\mathbb{R}^{2}$ such that $m\leq\inf_{\mathbb{R}}\alpha$ and $M\geq\sup_{\mathbb{R}}\alpha$. Then there exist two steady states of (3.12), $(v_{j})_{j}$ and $(w_{j})_{j}$, such that:
(i) $(v_{j})_{j}$ is bounded from below and for all $j\in\mathbb{Z}$, $v_{j}\leq m$ ;
(ii) $(w_{j})_{j}$ is bounded from above and for all $j\in\mathbb{Z}$, $w_{j}\geq M$.
Proof.  We only give the details for the construction of $(w_{j})_{j}$. Set
$\mathcal{U}_{h}\coloneq\max_{x\in\mathbb{R}}H(x,M)\quad\text{and}\quad\overline{M}\coloneq\mathcal{U}_{h}+\sup_{\underset{{\left|v\right|}\leq 1}{x\in\mathbb{R}}}L(x,v).$
By monotonicity, for all $j\in\mathbb{Z}$, $S_{j}^{+}(\mathcal{U}_{h})\geq M$. Therefore, taking advantage of Lemma 3.10 (ii):
$\forall j\in\mathbb{Z},\quad S_{j}^{+}(\mathcal{U}_{h})\leq H(x_{j},S_{j}^{+}(\mathcal{U}_{h}))+\sup_{\underset{{\left|v\right|}\leq 1}{x\in\mathbb{R}}}L(x,v)=\overline{M},$
ensuring that $\overline{M}\geq M\geq\sup_{\mathbb{R}}\alpha$. Now, define
$\forall j\in\mathbb{Z},\quad w_{j+1}\coloneq\left\{\begin{array}[]{ccl}\overline{M}&\text{if}&j\leq-J-2\\ S_{j+1}^{+}(h_{j}(w_{j}))&\text{if}&j\geq-J-1.\end{array}\right$.
Lemma 3.9 ensures that $(w_{j})_{j}$ is a steady state of the scheme since $\overline{M}\geq\sup_{\mathbb{R}}\alpha$.
@@
for all $j\in\mathbb{Z}$, $w_{j}\geq M$. The key point is that $(h_{j}(w_{j}))_{j}$ is constant. Therefore, for all $j\in\mathbb{Z}$,
$h_{j}(w_{j})=h_{-J-1}(w_{-J-1})=h_{-J-1}(\overline{M})\geq h_{-J-1}(S_{-J-1}^{+}(\mathcal{U}_{h}))=\mathcal{U}_{h}\geq h_{j}(M),$
from which we deduce, by monotonicity, that $w_{j}\geq M$. The claim is proved.
@@
$(w_{j})_{j}$ is bounded from above. Once again, we use that $(h_{j}(w_{j}))_{j}$ is constant. For all $j\in\mathbb{Z}$,
$h_{j}(w_{j})=h_{-J-1}(\overline{M})$ $\implies w_{j}=S_{j}^{+}(h_{-J-1}(\overline{M}))$
$\implies w_{j}=S_{j}^{+}\left(H(-X,\overline{M})\right)$
[…]
We see that $\overline{u}$ only depends on $H$ and $M$. Let us precise that one can choose
$\underline{u}\coloneq-H(-X,\overline{m})-\sup_{\underset{{\left|v\right|}\leq 1}{x\in\mathbb{R}}}L(x,v),\quad\overline{m}\coloneq-u_{h}-\sup_{\underset{{\left|v\right|}\leq 1}{x\in\mathbb{R}}}L(x,v),\quad u_{h}\coloneq\max_{x\in\mathbb{R}}H(x,m).$
$\square$
Proof of (M1).  With reference to Lemma 3.6, fix $(m,M)\in\mathbb{R}^{2}$ satisfying $m\leq\inf_{\mathbb{R}}\alpha$ and $M\geq\sup_{\mathbb{R}}\alpha$, such that for a.e. $x\in\mathbb{R}$, $m\leq u_{o}(x)\leq M$. Consider $(v_{j})_{j}$ and $(w_{j})_{j}$, steady states given by Lemma 3.11.
@@
$(v_{\Delta},w_{\Delta})\coloneq\sum_{j\in\mathbb{Z}}(v_{j},w_{j})\mathbf{1}_{I_{j}},$
and the constants $\underline{u}$, $\overline{u}$ as they were defined in the proof of Lemma 3.11.
The key point is that $v_{\Delta}$ and $w_{\Delta}$ are both exact stationary solutions to
$\partial_{t}\mathcal{U}(t,x)+\partial_{x}\bigl(H(\xi_{\Delta}(x),\mathcal{U}(t,x))\bigr)=0.$
Therefore, for all $n\in\mathbb{N}$, recalling the definition of $\mathcal{U}^{n}$, (D2) of Theorem 3.5 ensures that
$v_{\Delta}\leq u^{n}\leq w_{\Delta}\implies\forall t\in[t^{n},t^{n+1}],\quad v_{\Delta}\leq\mathcal{U}^{n}(t)\leq w_{\Delta}.$
Since by construction of $(v_{j})_{j\in\mathbb{Z}}$ and $(w_{j})_{j\in\mathbb{Z}}$, we have $v_{\Delta}\leq\rho^{0}\leq w_{\Delta}$, a direct induction provides:
$\forall t\in\mathbb{R}^{+},\quad v_{\Delta}\leq u_{\Delta}(t)\leq w_{\Delta},$
immediately leading to (2.4) by definition of $\underline{u}$ and $\overline{u}$. The proof is concluded.   $\square$
@@ 3.3 Convergence of the scheme
This section is devoted to the proof of the convergence of the scheme. First, we establish the strong compactness of the approximate solution using the compensated compactness method. Then, in Section 3.3, we prove that the scheme verifies discrete entropy inequalities and pass to the limit using the pre-established compactness.
Below, the constants $\underline{u},\overline{u}$ refer to the ones from (M1) of Theorem 2.1.